OpenStudy (anonymous):

Solve for t: 2t^2-20t-200=0

OpenStudy (anonymous):

Nice name there

OpenStudy (anonymous):

I know that it needs to be factored or put into the quadratic equation

OpenStudy (kirbykirby):

you can divide the equation by 2 on both sides: t^2 - 10t - 100 = 0 This just makes it easier to deal with smaller numbers. Since it doesn't appear you can factor it easily, you have to use the quadratic formula. Here, you have the form a+bx+cx^2 where a = 1, b=-10 and c=-10 so just use the quadratic equation

OpenStudy (kirbykirby):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} = \frac{-(-1)\pm \sqrt{(-)^2-4(1)(-100)}}{2(1)}\]

OpenStudy (kirbykirby):

sorry.. I messed up its -(-10) +/- sqrt (-10)^2.... the b = -10

OpenStudy (anonymous):

Thank you so much dude

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

That now explains it well